A rectangular piece of sheet metal has a length that is 4.3 times its width. In each corner a square with sides of 2.1 inches is cut out. The outer strips are then bent up to from an open box with a volume of 265 in^3.
Find the width (in inches) of the original sheet of metal.

Answer by Boreal(15235)

width is x
length is 4.3x
the new width after 2.1 in is cut from each side is x-4.2 and the new length is 4.3x-4.2
the height is 2.1 in
the volume is 265 in^3
Therefore, (4.3x^2-18.06x-4.2x+17.64)2.1=265
or 4.3x^2-18.06x-4.2x+17.64=126.19
4.3x^2-22.26x-108.55=0
x=(1/8.6)(22.26+/- sqrt (495.51+1867.06); sqrt=48.61
=(1/8.6)(22.26+48.61)=8.24 in width ANSWER
length is 35.43 in length
the box is 4.0 in x 31.2 x 2.1 =262.08 in ^3 (rounded)
4.04 in x 31.23 in x 2.1=264.96 in ^3